OpenStudy (anonymous):

need a number value for: 28^x=7

OpenStudy (anonymous):

28^x=7 ln28^x=ln7 xln28=ln7 x=ln7/ln28 put that in your calc

OpenStudy (anonymous):

Thanks I remember now!

OpenStudy (anonymous):

alright good!!